We assume the existence of a uniform electric field E= 500 i N/C without specifying which charges are the source. To this field is added that of a point charge of 2 uc located at the origin. What is the resultant electric force acting on a charge of 5 uc located at the point (3m,4m)?




The answer is (4.66i + 2.88j)x10^-3 N but I'm not sure how they got it?

1
Expert's answer
2022-02-17T04:36:07-0500

"F=qE+\frac{kqq_0}{r_x^2+r_y^2}=5.48~mN,"

"F_x=F\cos\alpha=4.66~mN,"

"F_y=F\sin\alpha=2.88~mN."